# Data-Driven Integration Kernels for Interpretable Nonlocal Operator Learning

### 11 Mar 2026 v1

@@ Abstract
Machine learning models can represent climate processes that are nonlocal in horizontal space, height, and time, often by combining information across these dimensions in highly nonlinear ways. While this can improve predictive skill, it makes learned relationships difficult to interpret and prone to overfitting as the extent of nonlocal information grows. We address this challenge by introducing data-driven integration kernels, a framework that adds structure to nonlocal operator learning by explicitly separating nonlocal information aggregation from local nonlinear prediction. Each spatiotemporal predictor field is first integrated using learnable kernels (defined as continuous weighting functions over horizontal space, height, and/or time), after which a local nonlinear mapping is applied only to the resulting kernel-integrated features and any optional local inputs. This design confines nonlinear interactions to a small set of integrated features and makes each kernel directly interpretable as a weighting pattern that reveals which horizontal locations, vertical levels, and past timesteps contribute most to the prediction. We demonstrate the framework for South Asian monsoon precipitation using a hierarchy of neural network models with increasing structure, including baseline, nonparametric kernel, and parametric kernel models. Across this hierarchy, kernel-based models achieve near-baseline performance with far fewer trainable parameters, showing that much of the relevant nonlocal information can be captured through a small set of interpretable integrations when appropriate structural constraints are imposed.
@@
nonlocal operator learning, interpretable machine learning, kernel integration
journal: Environmental Data Sciencearticletype: METHODS PAPER
@@ 1 Introduction
Geophysical processes are inherently nonlocal, with local outcomes depending on conditions across neighboring horizontal locations, the vertical column, and past timesteps. A growing body of machine learning research addresses this structure by learning operators [Kovachki2023] that map entire, high-dimensional input fields to local outputs. Such models have been used to learn nonlocal processes in weather and climate, and have achieved strong predictive performance across a range of applications [Pathak2022, Bonev2025, WattMeyer2025]. These operator-learning approaches can flexibly combine information across horizontal space, height, and time, but nonlocal contributions are typically encoded implicitly within large parameter sets, making it difficult to identify which spatial scales, vertical levels, or memory timescales are most influential. As a result, adding broader nonlocal context can increase model complexity without providing clearer insight into how nonlocal information is actually used.
A range of strategies have been proposed to manage this complexity. Dimensionality-reduction approaches such as principal-component analysis and encoder–decoder architectures (including autoencoders) compress high-dimensional fields into lower-dimensional latent representations, but these representations often reflect architectural or statistical choices rather than physically meaningful structure [Monahan2009, Wani2025]. Post-hoc explainability methods aim to attribute predictions to influential regions or features, but their results depend on both the trained model and the attribution technique [Silva2024, OLoughlin2025]. Because such interpretations are produced by fitting additional explanatory models after training, they can compound modeling uncertainty, reinforce prior expectations, and fail to provide stable or physically interpretable summaries of nonlocal influence.
Together, these limitations point to the need for approaches that represent nonlocal structure explicitly within the learned operator, while retaining the expressive power required to model complex geophysical systems. Such approaches should reduce dimensionality, make nonlocal influence directly inspectable, and support systematic comparison across models and data settings, without relying on post-hoc interpretation or latent variables whose meaning depends on model-specific choices.
@@ 2.1 Nonlocal Operators via Integration Kernels
[…]
@@ 3.2 Implementation Framework
The thermodynamic predictor fields are relative humidity (RH), equivalent potential temperature ($\theta_{e}$), and saturated equivalent potential temperature ($\theta_{e}^{*}$), which are relevant for South Asian monsoon rainfall variability [Anonymous2025]. Local inputs (evaluated only at the prediction point) are sensible heat flux, latent heat flux, and land fraction. ERA5 reanalysis [Hersbach2023a, Hersbach2023b] provides physically consistent, well-resolved vertical profiles over the South Asian monsoon region, particularly for lower-tropospheric temperature and moisture [Hersbach2020, Johnston2021]. Accordingly, all predictor fields and local inputs are extracted from hourly ERA5 data on its native $0.25^{\circ}\times 0.25^{\circ}$ grid. Thermodynamic predictors are constructed from surface pressure and from temperature and specific humidity on pressure levels between 1,000 and 500 hPa. Precipitation targets are taken from half-hourly IMERG V06 [Huffman2019] on its native $0.1^{\circ}\times 0.1^{\circ}$ grid, and are aggregated to hourly using centered hourly means. Prior to model training, both datasets are conservatively regridded to a common $1.0^{\circ}\times 1.0^{\circ}$ grid, consistent with the horizontal scales at which precipitation is typically parameterized in large-scale atmospheric models.
@@
We use a chronological split with data from 2000–2014 for training, 2015–2017 for validation, and 2018–2020 for testing. All variables except land fraction (since it is time-invariant and bounded in $[0,1]$) are standardized using training set means and standard deviations, with precipitation transformed using $\log(1+y)$ prior to standardization to reduce the influence of extremes and improve training stability.
Figure 1 provides a schematic overview of the kernel-based modeling framework implemented in this study.
@@ 3.3 Model Hierarchy
At the third level, we train parametric kernel models that further constrain the aggregation stage by restricting kernels to simple functional families (e.g., Gaussian, mixture-of-Gaussians, top-hat, or exponential; see Appendix A). As in the nonparametric case, predictor fields are integrated using kernels and mapped to precipitation through the same downstream network. We consider both models that apply a single kernel family across all predictors and models that allow predictor-specific choices of parametric families.
@@ 4 Results
We evaluate the integration kernel approach by (1) examining how predictive performance depends on the representation of nonlocal structure and (2) analyzing the vertical kernel structure recovered by the kernel models. All results are reported on the test set and evaluated at the common set of prediction points described in Section 3.2.
@@
Figure 2 summarizes predictive performance across the model hierarchy introduced in Section 3.3. The fully local and fully nonlocal baselines provide lower and upper bounds on skill for the prescribed nonlocal domain, while intermediate baseline models that introduce nonlocality along one dimension at a time form an ablation study of horizontal, vertical, and temporal contributions. Among these, the vertically nonlocal model lies closest to the fully nonlocal upper bound, whereas models that are nonlocal only in space or only in time remain closer to the local lower bound. This ordering is consistent with the dominant role of vertical thermodynamic structure in controlling convective precipitation, with horizontal and temporal context playing a secondary role at the scales considered here [Arakawa1974]. That the model hierarchy reflects this expected behavior provides reassurance that the learned representations are capturing physically meaningful sources of predictive skill, motivating our focus on kernel analyses in the vertical dimension. Test set $\mathrm{R}^{2}$ values and variability across random seeds are reported for all models in Table A1 (Appendix C).
Relative to the fully local baseline ($\mathrm{R}^{2}\approx 0.41$), introducing vertical nonlocality increases $\mathrm{R}^{2}$ to $\sim$0.53 (a gain of $\sim$0.12). Relative to this vertically nonlocal baseline, the nonparametric kernel model recovers approximately $75\%$ of this gain, while the best parametric kernel model recovers approximately $67\%$. Differences among parametric families are modest (approximately $7\times 10^{-3}$ in $\mathrm{R}^{2}$; Table A1 in Appendix C), indicating that predictive skill is insensitive to the parametric form once vertical nonlocality is enforced. Kernel-based models therefore achieve performance close to the vertically nonlocal baseline despite operating on substantially reduced input representations, with the remaining loss of skill in the parametric case reflecting the stronger functional constraints imposed on the integration stage.
@@
Figure 2: Test set $R^{2}$ for baseline (blue), nonparametric kernel (yellow), and parametric kernel (red) models. Values are computed in the standardized, log-transformed space. Model labels indicate the dimensions treated nonlocally, with a subscript 0 denoting locality. Superscript k denotes nonparametric kernels, while TH, EXP, G, and MG denote parametric kernel families (top-hat, exponential, Gaussian, and mixture-of-Gaussians; see Appendix A). The MIX kernel uses mixture-of-Gaussians kernels for RH and $\theta_{e}^{*}$ and an exponential kernel for $\theta_{e}$
@@
Figure 3 shows that the learned vertical kernels exhibit distinct, predictor-specific structure across the lower troposphere, consistent with known physical controls on South Asian monsoon convection [Anonymous2025, Parker2016, Menon2018, Gayatri2024, Ahmed2021]. For RH, the nonparametric kernels emphasize both near-surface levels (approximately 900–1,000 hPa) and the lower free troposphere (roughly 650–500 hPa), reflecting the combined roles of boundary-layer moisture supply and free-tropospheric humidity in regulating convective intensity and precipitation efficiency. The $\theta_{e}$ kernels exhibit broadly positive weighting through the lower troposphere together with a localized negative contribution near $\sim$600 hPa, indicating sensitivity to the contrast between boundary-layer parcel energy and lower free-tropospheric conditions, rather than to $\theta_{e}$ at any single level. The $\theta_{e}^{*}$ kernels show alternating positive and negative contributions in the lower free troposphere, indicating sensitivity to the vertical distribution of stability and entrainment-driven dilution that determines whether boundary-layer buoyancy can be maintained during ascent.
@@
Parametric kernel models retain these dominant vertical sensitivities while imposing substantially stronger structural constraints. In our best parametric model, RH and $\theta_{e}^{*}$ are represented using mixture-of-Gaussians kernels, while $\theta_{e}$ is represented using an exponential kernel; choices motivated by the qualitative structure of the nonparametric kernels described above. Together, these parameterizations express vertical dependence using a small number of broad modes that capture the primary boundary layer and lower free-tropospheric contrasts while suppressing fine-scale variability. As a result, higher-order oscillations present in the nonparametric kernels are smoothed or omitted, yet the remaining structure preserves the physically relevant vertical controls identified above. Kernels for all predictors and kernel variants are shown in Figure A1 (Appendix C).
@@
Figure 3: Learned vertical kernels for each predictor field from the two best-performing kernel-based models from Figure 2. Columns correspond to predictor variables. The top row shows nonparametric kernels and the bottom row shows kernels from the best parametric model, with row labels following the notation introduced in Figure 2. Solid lines indicate the mean kernel weight across training seeds (n = 3), with shading denoting $\pm 1$ standard deviation. Larger absolute normalized kernel magnitude indicates greater contribution of a given pressure level, with sign indicating positive or negative influence. For predictors where mixture-of-Gaussians kernels were used, the dashed blue lines depict the individual components and the solid red line depicts their sum
@@ 5 Conclusion
We introduced data-driven integration kernels as an interpretable framework for learning nonlocal operators by separating nonlocal linear aggregation from local nonlinear prediction. This structure constrains model complexity, reduces dimensionality, and yields kernels that are directly interpretable as weighting patterns over horizontal space, height, and/or time. When this method is applied to South Asian monsoon precipitation, we find that most of the predictive skill gained from nonlocal information arises from vertical structure. This is consistent with established understanding that vertical thermodynamic profiles exert primary control on the generation of deep convective rainfall at the scales considered here, and it provides reassurance that the framework captures physically meaningful sources of predictive skill. At the same time, a key methodological result is that kernel-based models that restrict aggregation to this dimension retain performance close to the strongest baselines while using substantially fewer input features, with parametric kernels providing additional regularization, albeit with some loss in skill. Beyond prediction, kernel-integrated features provide compact summaries of nonlocal influence that are well suited for symbolic regression, enabling systematic comparison across models and datasets and supporting the development of analytic, physically interpretable relationships from data-driven operators.
@@ Appendix A Parametric Kernel Families
$k^{(\mathrm{G})}(s)=\exp\!\left(-\frac{\|s-\mu\|^{2}}{2\sigma^{2}}\right),$ (A.1)
@@
Mixture-of-Gaussians Kernel. To allow limited multimodal or opposing structure, we consider mixtures of Gaussian kernels. Using the Gaussian form above, a two-component mixture is defined as:
$k^{(\mathrm{MG})}(s)=w_{1}\exp\!\left(-\frac{\|s-\mu_{1}\|^{2}}{2\sigma_{1}^{2}}\right)+w_{2}\exp\!\left(-\frac{\|s-\mu_{2}\|^{2}}{2\sigma_{2}^{2}}\right).$ (A.2)
where $\mu_{1},\mu_{2}$ are component centers, $\sigma_{1},\sigma_{2}>0$ are component widths, $\|\cdot\|$ denotes the Euclidean norm when $s=\mathbf{x}$, and $w_{1},w_{2}$ are learned scalar weights that may be positive or negative.
Top-Hat Kernel. The top-hat kernel represents uniform averaging over a bounded interval of the integration domain and is implemented using a smooth approximation for differentiability:
$k^{(\mathrm{TH})}(s)=\left(1+e^{-(s-a)/\epsilon}\right)^{-1}\left(1+e^{-(b-s)/\epsilon}\right)^{-1},$ (A.3)
where $a$ and $b$ are learned bounds, and $\epsilon>0$ controls the sharpness of the transition. In the limit $\epsilon\to 0$, this recovers uniform averaging over $\min(a,b),\,\max(a,b)]$.
@@
$k^{(\mathrm{EXP})}(s)=\exp\!\left(-\frac{\ell(s)}{\tau_{0}}\right),$ (A.4)
where $\tau_{0}>0$ is a learned decay scale, and $\ell(s)\geq 0$ measures distance to the anchor. For $s=\mathbf{x}$, distance is Euclidean, $\ell(\mathbf{x}_{n})=\|\mathbf{x}_{n}-\mathbf{x}_{0}\|$, where $\mathbf{x}_{0}$ is the prediction location and $\mathbf{x}_{n}$ is a location in the horizontal neighborhood. For $s=p$, distance is defined on the discrete pressure grid. Let $j\in\{0,\dots,N_{p}-1\}$ index pressure levels from top ($j=0$, 500 hPa) to bottom ($j=N_{p}-1$, 1,000 hPa), and define $\ell(p_{j})=(1-\alpha)\,j+\alpha\,(N_{p}-1-j)$, where $\alpha\in(0,1)$ is a learned parameter controlling whether decay is anchored near the top of the column ($\alpha\approx 0$) or near the surface ($\alpha\approx 1$). For $s=t$, distance is defined on the discrete time grid. Let $j\in\{0,\dots,N_{t}-1\}$ index timesteps from the start of the window ($j=0$ at $t_{0}-\tau_{\max}$) to the prediction time ($j=N_{t}-1$ at $t_{0}$), and define $\ell(t_{j})=(N_{t}-1)-j$, so influence decays exponentially backward into the past.
For numerical stability during optimization, small epsilon values ($10^{-8}$) are added to top-hat and mixture-of-Gaussians kernels to prevent exact zeros, and exponential decay scales $\tau_{0}$ are clamped to the range $10^{-4},100]$.
Different parametric kernel families may be assigned to different predictor fields within the same model, allowing the nonlocal aggregation stage to reflect predictor-specific structure while preserving a shared downstream mapping and a common set of kernel families.

### 31 Mar 2026 v2

@@ Abstract
Machine learning models can represent climate processes that are nonlocal in horizontal space, height, and time, often by combining information across these dimensions in highly nonlinear ways. While this can improve predictive skill, it makes learned relationships difficult to interpret and prone to overfitting as the extent of nonlocal information grows. We address this challenge by introducing data-driven integration kernels, a framework that adds structure to nonlocal operator learning by explicitly separating nonlocal information aggregation from local nonlinear prediction. Each spatiotemporal predictor field is first integrated using learnable kernels (defined as continuous weighting functions over horizontal space, height, and/or time), after which a local nonlinear mapping is applied only to the resulting kernel-integrated features and optional local inputs. This design confines nonlinear interactions to a small set of integrated features and makes each kernel directly interpretable as a weighting pattern that reveals which horizontal locations, vertical levels, and past timesteps contribute most to the prediction. We demonstrate the framework for South Asian monsoon precipitation using a hierarchy of neural network models with increasing structure, including baseline, nonparametric kernel, and parametric kernel models. Across this hierarchy, kernel models achieve near-baseline performance with far fewer trainable parameters, indicating that much of the relevant nonlocal information can be captured through a small set of interpretable integrations when appropriate structural constraints are imposed.
@@
nonlocal operator learning; interpretable machine learning; kernel integration
articletype: METHODS PAPERjournal: Environmental Data Science
@@ 1 Introduction
Geophysical processes are inherently nonlocal, with local outcomes depending on conditions across neighboring horizontal locations, the vertical column, and past timesteps. A growing body of machine learning research addresses this structure by learning operators [Kovachki2024] that map entire, high-dimensional input fields to local outputs. Such models have been used to learn nonlocal processes in weather and climate, and have achieved strong predictive performance across a range of applications [Pathak2022, Bonev2025, WattMeyer2025]. These operator-learning approaches can flexibly combine information across horizontal space, height, and time, but nonlocal contributions are typically encoded implicitly within large parameter sets, making it difficult to identify which spatial scales, vertical levels, or memory timescales are most influential. As a result, adding broader nonlocal context can increase model complexity without providing clearer insight into how nonlocal information is actually used.
A range of strategies have been proposed to manage this complexity. Dimensionality-reduction approaches such as principal-component analysis and encoder–decoder architectures (including autoencoders) compress high-dimensional fields into lower-dimensional latent representations, but these representations often reflect architectural or statistical choices rather than physically meaningful structure [King2025, Monahan2009, Wani2025]. Post-hoc explainability methods aim to attribute predictions to influential regions or features, but depend on both the trained model and the attribution technique. As a result, they can mix contributions across predictors and dimensions and yield sample-dependent attributions that are difficult to interpret physically [Silva2024, OLoughlin2025]. Because these approaches rely on additional models fitted after training, they can compound uncertainty, reinforce prior expectations, and fail to provide stable summaries of nonlocal influence.
@@ 2.1 Nonlocal Operators via Integration Kernels
Formally, the $\ell^{\mathrm{th}}$ kernel-integrated feature associated with predictor $\varphi_{i}$ is defined as:
@@ 3.2 Implementation Framework
The thermodynamic predictor fields are relative humidity (RH), equivalent potential temperature ($\theta_{e}$), and saturated equivalent potential temperature ($\theta_{e}^{*}$), which are relevant for South Asian monsoon rainfall variability [Ferretti2025]. Local inputs (evaluated only at the prediction point) are sensible heat flux, latent heat flux, and land fraction. ERA5 reanalysis [Hersbach2023a, Hersbach2023b] provides physically consistent, well-resolved vertical profiles over the South Asian monsoon region, particularly for lower-tropospheric temperature and moisture [Hersbach2020, Johnston2021]. Accordingly, all predictor fields and local inputs are extracted from hourly ERA5 data on its native $0.25^{\circ}\times 0.25^{\circ}$ grid. Thermodynamic predictors are constructed from surface pressure and from temperature and specific humidity on pressure levels between 1,000 and 500 hPa. Precipitation targets are taken from half-hourly IMERG V06 [Huffman2019] on its native $0.1^{\circ}\times 0.1^{\circ}$ grid, and are aggregated to hourly resolution using centered averages. Prior to model training, both datasets are conservatively regridded to a common $1.0^{\circ}\times 1.0^{\circ}$ grid, consistent with the horizontal scales at which precipitation is typically parameterized in large-scale atmospheric models.
@@
We use a chronological split with data from 2000–2014 for training, 2015–2017 for validation, and 2018–2020 for testing. All variables except land fraction (since it is time-invariant and bounded in $[0,1]$) are standardized using training set means and standard deviations, with precipitation log1p-transformed prior to standardization to reduce the influence of extremes and improve training stability.
Figure 1 provides a schematic overview of the kernel modeling framework implemented in this study.
@@ 3.3 Model Hierarchy
At the third level, we train parametric kernel models that further constrain the aggregation stage by restricting kernels to simple functional families (e.g., Gaussian, mixture-of-Gaussians, top-hat, or exponential; see Appendix A). As in the nonparametric case, predictor fields are integrated using these kernels and then mapped to precipitation through the same downstream network. We consider both configurations that apply a single kernel family across all predictors and those that allow predictor-specific choices. In this sense, the selected parametric families are informed by the qualitative structure of the nonparametric kernels for this application, but the framework itself does not impose these choices, and alternative functional forms can be adopted to better match different datasets or physical regimes.
@@ 4 Results
We evaluate the integration kernel approach by (1) examining how predictive performance depends on the representation of nonlocal structure and (2) analyzing the vertical kernel structure recovered by the kernel models. All results are reported on the test set at the common prediction points defined in Section 3.2.
Figure 2 summarizes predictive performance across the model hierarchy introduced in Section 3.3, using test set R2 and mean squared error (MSE). R2 provides a scale-independent measure of explained variance, while MSE quantifies absolute error aligned with the training objective. Table A1 (see Appendix C) reports corresponding uncertainties from bootstrap resampling and variability across random seeds, which are uniformly small ($\mathcal{O}(10^{-3})$), indicating robust performance differences across sampling and initialization.
@@
Considering only the baseline models in Figure 2, the fully local $(\mathbf{x}_{0},p_{0},t_{0})$ and fully nonlocal $(\mathbf{x},p,t)$ models define lower and upper bounds on skill for the prescribed nonlocal domain (R2 = 0.411, MSE = 0.680 and R2 = 0.582, MSE = 0.482, respectively). Intermediate models that introduce nonlocality along a single dimension provide an ablation of horizontal, vertical, and temporal contributions. Among these, the vertically nonlocal model (R2 = 0.528, MSE = 0.546) lies closest to the upper bound, whereas the horizontally (R2 = 0.440, MSE = 0.647) and temporally (R2 = 0.432, MSE = 0.656) nonlocal models remain closer to the lower bound. This ordering is consistent with the dominant role of vertical thermodynamic structure in controlling convective precipitation, with horizontal and temporal context playing a secondary role at the scales considered here [Arakawa1974]. That this hierarchy within the baseline models reflects expected physical behavior provides reassurance that the learned representations capture physically meaningful sources of predictive skill, motivating our focus on developing kernels in the vertical dimension.
Relative to the vertically nonlocal baseline model, kernel models in Figure 2 show that the nonparametric kernel model achieves comparable performance despite operating on substantially reduced input representations (R2 = 0.496 versus 0.528, MSE = 0.582 versus 0.546). The parametric kernel models perform slightly worse than the nonparametric kernel model (R2 $\approx$ 0.481–0.488, MSE $\approx$ 0.592–0.599), with only minor spread across formulations, indicating limited sensitivity to the specific parametric form once vertical nonlocality is enforced. The remaining degradation of the parametric kernel models relative to the nonparametric kernel model reflects stronger functional constraints in the integration stage. Spatial composites for dry, typical, and wet conditions show consistent qualitative behavior across the vertically nonlocal baseline model, the best nonparametric kernel model, and the best parametric kernel model (Figure A1, see Appendix C), demonstrating that these models reproduce ground truth patterns closely for various precipitation regimes.
@@
Figure 2: Test set R2 (top) and MSE (bottom) for baseline (blue), nonparametric kernel (yellow), and parametric kernel (red) models, computed in standardized log1p-transformed precipitation space. Model labels indicate nonlocal dimensions, with subscript 0 denoting locality. Superscript $k$ denotes nonparametric kernels, while TH, EXP, G, and MG denote top-hat, exponential, Gaussian, and mixture-of-Gaussians kernels (see Appendix A). MIX uses MG kernels for RH and $\theta_{e}^{*}$ and EXP for $\theta_{e}$
@@
Figure 3 shows that the learned vertical kernels exhibit distinct, predictor-specific structure across the lower troposphere, consistent with known physical controls on South Asian monsoon convection [Ferretti2025, Parker2016, Menon2018, Gayatri2024, Ahmed2021]. For RH, the nonparametric kernels emphasize both near-surface levels (approximately 900–1,000 hPa) and the lower free troposphere (roughly 650–500 hPa), reflecting the combined roles of boundary-layer moisture supply and free-tropospheric humidity in regulating convective intensity and precipitation efficiency. The $\theta_{e}$ kernels exhibit broadly positive weighting through the lower troposphere, with a localized negative contribution near 600 hPa, indicating sensitivity to the contrast between boundary-layer parcel energy and lower free-tropospheric conditions, rather than to $\theta_{e}$ at any single level. The $\theta_{e}^{*}$ kernels show alternating positive and negative contributions in the lower free troposphere, indicating sensitivity to the vertical distribution of stability and entrainment-driven dilution that determines whether boundary-layer buoyancy can be maintained during ascent [Ahmed2018].
@@
Parametric kernel models retain these dominant vertical sensitivities while imposing stronger structural constraints. In our best parametric model, RH and $\theta_{e}^{*}$ are represented using mixture-of-Gaussians kernels, while $\theta_{e}$ is represented using an exponential kernel, choices motivated by the qualitative structure of the nonparametric kernels described above. Together, these parameterizations capture the primary boundary-layer and lower free-tropospheric sensitivities while suppressing fine-scale variability, with higher-order oscillations and sign-changing structure absorbed into the downstream nonlinear mapping. The parametric $\theta_{e}^{*}$ kernel in particular collapses onto a dominant positive peak near 600 hPa, where the moisture deficit ($\theta_{e}^{*}-\theta_{e}$) becomes increasingly important for regulating entrainment-driven dilution of rising parcels, consistent with the positive RH weighting at similar levels. That the parametric and nonparametric $\theta_{e}^{*}$ kernels differ most in shape reflects both the structural constraints of the mixture-of-Gaussians kernel and underdetermination, since multiple kernel profiles can achieve similar predictive performance. Kernels for all predictors and kernel models are shown in Figure A2 (see Appendix C).
@@
Figure 3: Learned vertical kernels for predictor fields from the two best kernel models in Figure 2. Solid lines show mean weights across seeds ($n$ = 3), shaded by standard deviation. Absolute magnitude indicates pressure level importance, with sign indicating influence direction. For mixture-of-Gaussians cases, dashed blue lines show individual components and the solid red line shows their sum
@@ 5 Conclusion
We introduced data-driven integration kernels as an interpretable framework for learning nonlocal operators by separating nonlocal linear aggregation from local nonlinear prediction. This structure constrains model complexity, reduces dimensionality, and yields kernels that are directly interpretable as weighting patterns over horizontal space, height, and/or time. When applied to South Asian monsoon precipitation, we find that most of the predictive skill gained from nonlocal information arises from vertical structure. This is consistent with established understanding that vertical thermodynamic profiles exert primary control on the generation of deep convective rainfall at the scales considered here, and it provides reassurance that the framework captures physically meaningful sources of predictive skill. At the same time, a key methodological result is that kernel models that restrict aggregation to this dimension retain performance close to the strongest baselines while using substantially fewer input features, with parametric kernels providing additional regularization, albeit with some loss in skill. Beyond prediction, kernel-integrated features provide compact summaries of nonlocal influence that are well suited for symbolic regression, offering a potential route toward simpler analytic approximations of the learned operators. This may enable future work to more directly compare and interpret nonlocal mechanisms and assess their variation across weather regimes.
@@ Appendix A Parametric Kernel Families
$k^{(\mathrm{G})}(s)=\exp\!\left(-\tfrac{\lVert s-\mu\rVert^{2}}{2\sigma^{2}}\right),$ (A.1)
@@
Mixture-of-Gaussians Kernel. To allow bimodal or antagonistic structure, we consider mixtures of Gaussian kernels with learned scalar weights $w_{1},w_{2}$ that may be positive or negative. Negative weights enable one component to suppress the influence of another. A two-component mixture is defined as:
$k^{(\mathrm{MG})}(s)=w_{1}\exp\!\left(-\tfrac{\lVert s-\mu_{1}\rVert^{2}}{2\sigma_{1}^{2}}\right)+w_{2}\exp\!\left(-\tfrac{\lVert s-\mu_{2}\rVert^{2}}{2\sigma_{2}^{2}}\right),$ (A.2)
where $\mu_{1},\mu_{2}$ are component centers, $\sigma_{1},\sigma_{2}\in[0.1,2.0]$ are component widths (clamped for numerical stability), and $\|\cdot\|$ denotes the Euclidean norm when $s=\mathbf{x}$.
Top-Hat Kernel. The top-hat kernel represents uniform averaging over a bounded interval of the integration domain. It is implemented as a product of two sigmoid functions for differentiability:
$k^{(\mathrm{TH})}(s)=\Bigl(1+\exp\!\bigl(-\tfrac{s-a}{\epsilon}\bigr)\Bigr)^{-1}\Bigl(1+\exp\!\bigl(-\tfrac{b-s}{\epsilon}\bigr)\Bigr)^{-1},$ (A.3)
where $\epsilon=0.02$ is a fixed sharpness parameter, $a=\min(\ell,u)$ and $b=a+\tilde{w}$ are the effective lower and upper bounds, and $\ell,u$ are learned parameters. The effective width $\tilde{w}$ is derived from the raw learned width $w=\max(\ell,u)-\min(\ell,u)$ via soft clamping: $\tilde{w}=w$ for $w\leq 1.5$, and $\tilde{w}=1.5\tanh(w/1.5)$ otherwise, preventing the kernel from spanning the full normalized domain.
@@
$k^{(\mathrm{EXP})}(s)=\exp\!\left(-\tfrac{d(s)}{\tau_{0}}\right),$ (A.4)
where $\tau_{0}>0$ is a learned decay scale and $d(s)\geq 0$ measures distance to the anchor. For $s=\mathbf{x}$, distance is Euclidean, $d(\mathbf{x}_{n})=\|\mathbf{x}_{n}-\mathbf{x}_{0}\|$, where $\mathbf{x}_{0}$ is the prediction location and $\mathbf{x}_{n}$ is a location in the horizontal neighborhood. For $s=p$, distance is defined on the discrete pressure grid. Let $j\in\{0,\dots,N_{p}-1\}$ index pressure levels from top ($j=0$, 500 hPa) to bottom ($j=N_{p}-1$, 1,000 hPa), and define $d(p_{j})=(1-\alpha)\,j+\alpha\,(N_{p}-1-j)$, where $\alpha\in(0,1)$ is a learned parameter. This is a linear interpolation between $j$ (index distance from the top) and $N_{p}-1-j$ (index distance from the bottom), so $\alpha\approx 0$ anchors decay near the top of the column and $\alpha\approx 1$ anchors it near the surface. For $s=t$, distance is defined on the discrete time grid. Let $j\in\{0,\dots,N_{t}-1\}$ index timesteps from the start of the window ($j=0$ at $t_{0}-\tau_{\max}$) to the prediction time ($j=N_{t}-1$ at $t_{0}$), and define $d(t_{j})=(N_{t}-1)-j$, the number of timesteps before the prediction time, so influence decays exponentially backward into the past.
For numerical stability during optimization, small values ($10^{-8}$) are added to top-hat and mixture-of-Gaussians kernels to prevent exact zeros, and exponential decay scales $\tau_{0}$ are clamped to the range $10^{-4},100]$. Different parametric kernel families may be assigned to different predictor fields within the same model, allowing the nonlocal aggregation stage to reflect predictor-specific structure while preserving a shared downstream mapping and a common set of kernel families.